# Near-extremal hydrodynamics and the holographic product formula

### v1

@@ 2.3 Special behavior of the soft poles
where $\omega^{\text{IR}}_{n}$ are the poles of the IR CFTp-1 correlator $\mathcal{G}_{12}$. Examples of this behavior for $p=2$ are given in Edalati:2010hk; Berkooz:2012qh; Davison:2022vqh; Jia:2024zes; Arnaudo:2024sen (with Gouteraux:2025kta being a counter-example), whereas section 3.3 describes an example with $p=3$. Based on known examples, it looks like (20) typically holds either for correlators that do not have gapless poles (e.g. the transverse current in section 3.1 or the scalar in section 3.3), or that couple to energy and momentum (e.g. the correlators of section 3.2).
@@
which is valid for all $\omega$ outside small disks of radius of order $\mathcal{O}(r_{e}T)$ around the IR poles. The generalized susceptibility $\tilde{\gamma}_{e}(\omega,\vec{k})$ is analytic in $\omega$ in the near-extremal hydrodynamic regime (11).
In the general near-extremal regime (11), $\mathcal{G}_{12}(\omega,\vec{k})$ is the two sided correlator in the IR CFTp-1 at finite temperature. This makes equation (23) particularly useful in the case of $p=2$ and 3, since the general expression of thermal two-point functions is known for one-dimensional and two-dimensional CFT’s (see e.g. Faulkner:2009wj and Dodelson:2023vrw; Son:2002sd). In these cases, the soft factor (22) can therefore be expressed at leading order in $r_{e}T$ in an exact analytic form (up to a $\vec{k}$-dependent factor), including for energies $|\omega|\lesssim T$.
@@ 3.1 Probe current correlators
The operator dual to $A_{\mu}$ is a conserved current $J^{\mu}$, with vanishing background density $\left<J^{0}\right>=0$, and not charged under any other background density. This corresponds to the correlators studied in Jarvinen:2023xrx, that are related to neutrino transport.
@@
where $d\geq 3$ is the boundary dimension [^12] We focus on theories that feature diffusion, which excludes in particular (1+1)-dimensional CFT’s Kovtun:2008kx; Herzog:2007ij. and the transverse gauge field and longitudinal electric field are defined as
@@
This potential has a pole when $f(z)\vec{k}^{2}=\omega^{2}$, which can be reached at finite $z$ for $\omega^{2}/\vec{k}^{2}<1$. Also its UV asymptotics are such that $\nu^{2}=(d-2)^{2}/4$, except when $\omega^{2}=\vec{k}^{2}$, for which $\nu^{2}=-3$. As a consequence, we expect the corresponding two-sided correlator to feature some zeroes for $\omega^{2}/\vec{k}^{2}\leq 1$. It can be checked (see e.g. Jarvinen:2023xrx) that the correlator associated with $E^{\parallel}$ is the polarization function $\Pi^{\parallel}$ in (26). In this case, it is therefore simple to understand why there is a zero: since the full correlator should not have a pole at $\omega^{2}=\vec{k}^{2}$, and because $\Pi^{\parallel}$ is multiplied by the longitudinal projector (27), it results that $\Pi^{\parallel}$ should vanish when $\omega^{2}=\vec{k}^{2}$.
@@
At leading order in the near-extremal hydrodynamic expansion, there is a single gapless pole $\omega_{D}(\vec{k})=-iD\vec{k}^{2}$, with $D=r_{H}/(d-2)$ the diffusion constant Davison:2013bxa. In this regime, the longitudinal spectral function may thus be written as
@@
with $\sigma$ the conductivity. The expression (42) corresponds to the standard hydrodynamic expression for a diffusive mode Jarvinen:2023xrx, improved with the appropriate IR power-law.
At next-to-leading order, the numerical analysis presented in appendix D indicates that $\omega_{D}$ splits into two hydrodynamic poles, much like in Gouteraux:2025kta (see (138)). At this order and higher, we therefore expect the polarization function to be described by an expression of the form (139) in the extremal hydrodynamic regime.
@@
with $\gamma$ analytic and free of zeroes, and $\mathcal{G}^{s}_{12}$ the soft factor (15). A numerical analysis of the transverse fluctuation equation (33) further indicates that the scenario discussed in section 2.3 applies in this sector, with the soft poles approaching the IR poles in the limit $r_{e}T\to 0$ (see appendix D). The near-extremal transverse polarization function may thus be written in the form (23), i.e. an analytic factor times the IR AdS2 correlator Faulkner:2009wj
@@ 3.2 Stress-tensor correlators in the sound channel
We now consider the case of helicity-0 stress-tensor correlators on the Reissner-Nordtröm (RN) background, whose gapless sector is qualitatively different from the previous example, since it features two types of gapless modes: a diffusive-like mode $\omega_{D}(\vec{k})=-iD\vec{k}^{2}+\mathcal{O}(\vec{k}^{4})$, and a sound-like mode $\omega_{s}(\vec{k})=c_{s}|\vec{k}|-i\Gamma\vec{k}^{2}+\mathcal{O}(\vec{k}^{4})$ Edalati:2010pn; Davison:2011uk; Moitra:2020dal.
@@
The stress-tensor correlators are computed from the fluctuation equations for the appropriate modes of the metric $g_{MN}$, which couple to the fluctuations of the gauge field (50). These equations can be written in terms of two gauge-invariants Kodama:2003kk; Edalati:2010pn
@@
The fields $\mathcal{A}$ and $\Phi$ obey Kodama:2003kk; Edalati:2010pn
@@
It is possible to further simplify the problem by introducing the following “master” variables [^14] Note the relations between our variables and those of Kodama-Ishibashi (KI) Kodama:2003kk $\mathcal{A}=\vec{k}^{2}\mathcal{A}_{\text{KI}}\quad,\quad\Phi=\frac{1}{n}\vec{k}^{2}\Phi_{\text{KI}}\,$, $\Phi_{+}=\frac{\vec{k}^{2}}{4nQ}\Phi_{-,\text{KI}}\quad,\quad\Phi_{-}=-\frac{\vec{k}^{2}}{(n+1)(M+\rho)}\Phi_{+,\text{KI}}\,.$
@@
We would now want to relate the stress-tensor and current correlators to the master fields Green’s functions $G^{\pm}(\omega,\vec{k})$. This relation depends on the dimensionality $n$, so we will now consider a fixed value of $n$. The case of $n=2$ was treated in Edalati:2010pn, where the sound channel correlators were found to be non-linearly related to $G^{\pm}$, as reviewed in appendix F. Here we will instead consider $n=3$, for which the relation becomes linear. We argue in appendix F that this also holds for any $n\geq 4$.
@@
From the definitions (56), and using the near-boundary equations of motion [^16] The latter imply in particular the Ward identities obeyed by the stress-tensor and current correlators (68)-(70). See Edalati:2010pn for more details of this calculation in the case $n=2$., the coefficients $\Phi_{\pm}^{(0)}$ and $\Phi_{\pm}^{(1)}$ can now be related to the metric and gauge field expansions (66)-(67) as
@@
Now taking the imaginary parts on both sides of equations (77)-(82) - and remembering that the contact terms are real - the sound channel two-sided correlators are found to be linearly related to the master correlators $G^{\pm}_{12}$. Since the latter obey the product formula, the stress tensor and current correlators can therefore also be written as a product over their poles. In particular, the $G^{-}$ correlator has a diffusive pole $\omega_{D}(k)=-iD\vec{k}^{2}+\mathcal{O}(\vec{k}^{4})$, whereas $G^{+}$ features a sound pole $\omega_{s}(\vec{k})=c_{s}|\vec{k}|-i\Gamma\vec{k}^{2}+\mathcal{O}(\vec{k}^{4})$ Arean:2020eus, as can be seen explicitly in the numerical results presented in appendix G. These results also indicate that the soft poles approach the IR poles in the near-extremal limit, as in (20). Following the discussion of section 2.3, in the near-extremal hydrodynamic regime we may therefore write the $G^{\pm}$ correlators as
$G^{-}_{12}(\omega,\vec{k})=\frac{\gamma_{-}(\omega,\vec{k})\mathcal{G}^{-}_{12}(\omega,\vec{k})}{\omega^{2}-\omega_{D}(\vec{k})^{2}}\,$, (87)
$G^{+}_{12}(\omega,\vec{k})=\frac{\gamma_{+}(\omega,\vec{k})\mathcal{G}^{+}_{12}(\omega,\vec{k})}{(\omega^{2}-\omega_{s}(\vec{k})^{2})(\omega^{2}-\omega_{s}^{*}(\vec{k})^{2})}\,$, (88)
with $\gamma_{\pm}(\omega,\vec{k})$ analytic in $\omega$ and free of zeroes (except potentially for $\vec{k}^{2}<0$), and $\mathcal{G}^{\pm}_{12}(\omega,\vec{k})$ the IR CFT1 correlators (44), with conformal dimensions
@@
$\text{Im}G_{\text{L}}(\omega,\vec{k})=\frac{\gamma_{L}(\omega,\vec{k})\text{Im}\mathcal{G}^{+}_{R}(\omega,\vec{k})}{(\omega^{2}-\omega_{D}(\vec{k})^{2})(\omega^{2}-\omega_{s}(\vec{k})^{2})(\omega^{2}-\omega_{s}^{*}(\vec{k})^{2})}\,,\quad\text{L}\in\{T,J,\text{mix}\}\,$, (90)
@@
  • As expected Moitra:2020dal, the leading order extremal hydrodynamic expressions
@@
    agree with the extrapolation of leading order hydrodynamics to zero temperature Kovtun:2012rj, up to the modification of the low frequency power-law behavior from $\omega$ to $\omega^{2\Delta_{+}(\vec{k})-1}$;
@@
  • For the energy-energy, density-density and energy-density spectral functions, the height of the sound peak is of order $\mathcal{O}((r_{e}|\vec{k}|)^{2\Delta_{+}(\vec{k})-3})$, whereas the diffusive peak is of order $\mathcal{O}((r_{e}|\vec{k}|)^{4\Delta_{+}(\vec{k})-2})$. This indicates that the diffusive peak is suppressed at low temperatures, as was observed in Davison:2011uk for $n=2$, and further provides the precise momentum scaling of this suppression in the extremal hydrodynamic regime. In particular, this scaling indicates that the diffusive peak tends to become more important as the momentum is increased, as was also observed in Davison:2011uk.
@@ 3.3 Magnetized branes
For our last example, we consider a black brane background with $p=3$, meaning that in the near-extremal regime the theory flows to a CFT2 in the IR. A solution of this sort was found in DHoker:2009mmn, corresponding to a magnetized black brane at zero density. The action for the system is Einstein-Maxwell
@@
The near-extremal magnetic brane solution DHoker:2009mmn interpolates between the BTZ solution (101) near the horizon, and $AdS_{d+1}$ in the UV. For the fields of the ansatz (97), this means that when the boundary is approached ($r\to 0$) the fields behave as
$A(r)=-\log{\left(\frac{r}{\ell}\right)}+\mathcal{O}(r^{4})\quad,\quad f(r)=1+\mathcal{O}(r^{4})\quad,\quad h(r)=1+\mathcal{O}(r^{4})\,$, (104)
@@
The potential (113) is regular everywhere and has UV asymptotics of the form (8) with $\nu^{2}=(d-1)^{2}/4\geq 0$, which ensures that the scalar two-sided correlator $G_{12}(\omega,k_{x},k_{y})$ associated with $\phi$ obeys the product formula without zeroes. The poles of the retarded correlator associated with $\phi$ can be computed from a numerical calculation, whose results are presented in appendix H. These results indicate that $\phi$ does not feature any gapless pole, and its soft poles approach the IR CFT2 poles at low temperature. According to the discussion of section (2.3), the near-extremal limit of $G_{12}$ may then be written as
$G_{12}(\omega,k_{x},k_{y})=\gamma_{e}(\omega,k_{x},k_{y})\mathcal{G}^{k_{y}}_{12}(\omega,k_{x})\,$, (114)
@@
The zero-temperature effective susceptibility $\tilde{\gamma}_{e}$ is a function of $r_{e}^{2}(\omega^{2}-k_{x}^{2})$ and $r_{e}^{2}k_{y}^{2}$, which is free of zeroes and analytic in its first argument.

### v2

@@ 2.3 Special behavior of the soft poles
where $\omega^{\text{IR}}_{n}$ are the poles of the IR CFTp-1 correlator $\mathcal{G}_{12}$. Examples of this behavior for $p=2$ are given in Edalati et al. (2010b); Berkooz et al. (2012); Davison et al. (2023); Jia and Rangamani (2024); Arnaudo and Withers (2025) (with Goutéraux et al. (2025) being a counter-example), whereas section 3.3 describes an example with $p=3$. Based on known examples, it looks like (20) typically holds either for correlators that do not have gapless poles (e.g. the transverse current in section 3.1 or the scalar in section 3.3), or that couple to energy and momentum (e.g. the correlators of section 3.2).
@@
which is valid for all $\omega$ outside small disks of radius of order $\mathcal{O}(r_{e}T)$ around the IR poles. The generalized susceptibility $\tilde{\gamma}_{e}(\omega,\vec{k})$ is analytic in the near-extremal hydrodynamic regime (11).
In the general near-extremal regime (11), $\mathcal{G}_{12}(\omega,\vec{k})$ is the two sided correlator in the IR CFTp-1 at finite temperature. This makes equation (23) particularly useful in the case of $p=2$ and 3, since the general expression of thermal two-point functions is known for one-dimensional and two-dimensional CFT’s (see e.g. Faulkner et al. (2011) and Dodelson et al. (2024); Son and Starinets (2002)). In these cases, the soft factor (22) can therefore be expressed at leading order in $r_{e}T$ in a closed form (up to a $\vec{k}$-dependent factor), including for energies $|\omega|\lesssim T$.
@@ 3.1 Probe current correlators
The operator dual to $A_{\mu}$ is a conserved current $J^{\mu}$, with vanishing background density $\left<J^{0}\right>=0$, and not charged under any other background density. This corresponds to the correlators studied in Järvinen et al. (2023), that are related to neutrino transport.
@@
where $d\geq 3$ is the boundary dimension [^12] We focus on theories that feature diffusion, which excludes in particular (1+1)-dimensional CFT’s Kovtun and Ritz (2008); Herzog et al. (2007). and the transverse gauge field and longitudinal electric field are defined as
@@
This potential has a pole when $f(z)\vec{k}^{2}=\omega^{2}$, which can be reached at finite $z$ for $\omega^{2}/\vec{k}^{2}<1$. Also its UV asymptotics are such that $\nu^{2}=(d-2)^{2}/4$, except when $\omega^{2}=\vec{k}^{2}$, for which $\nu^{2}=-3$. As a consequence, we expect the corresponding two-sided correlator to feature some zeroes for $\omega^{2}/\vec{k}^{2}\leq 1$. It can be checked (see e.g. Järvinen et al. (2023)) that the correlator associated with $E^{\parallel}$ is the polarization function $\Pi^{\parallel}$ in (26). In this case, it is therefore simple to understand why there is a zero: since the full correlator should not have a pole at $\omega^{2}=\vec{k}^{2}$, and because $\Pi^{\parallel}$ is multiplied by the longitudinal projector (27), it results that $\Pi^{\parallel}$ should vanish when $\omega^{2}=\vec{k}^{2}$.
@@
At leading order in the near-extremal hydrodynamic expansion, there is a single gapless pole $\omega_{D}(\vec{k})=-iD\vec{k}^{2}$, with $D=r_{H}/(d-2)$ the diffusion constant Davison and Parnachev (2013). In this regime, the longitudinal spectral function may thus be written as
@@
with $\sigma$ the conductivity. The expression (42) corresponds to the standard hydrodynamic expression for a diffusive mode Järvinen et al. (2023), improved with the appropriate IR power-law.
At next-to-leading order, the numerical analysis presented in appendix D indicates that $\omega_{D}$ splits into two hydrodynamic poles, much like in Goutéraux et al. (2025) (see (138)). At this order and higher, we therefore expect the polarization function to be described by an expression of the form (139) in the extremal hydrodynamic regime.
@@
with $\gamma$ analytic and free of zeroes, and $\mathcal{G}^{s}_{12}$ the soft factor (15). A numerical analysis of the transverse fluctuation equation (33) further indicates that the scenario discussed in section 2.3 applies in this sector, with the soft poles approaching the IR poles in the limit $r_{e}T\to 0$ (see appendix D). The near-extremal transverse polarization function may thus be written in the form (23), i.e. an analytic factor times the IR AdS2 correlator Faulkner et al. (2011)
@@ 3.2 Stress-tensor correlators in the sound channel
We now consider the case of helicity-0 stress-tensor correlators on the Reissner-Nordtröm (RN) background, whose gapless sector is qualitatively different from the previous example, since it features two types of gapless modes: a diffusive-like mode $\omega_{D}(\vec{k})=-iD\vec{k}^{2}+\mathcal{O}(\vec{k}^{4})$, and a sound-like mode $\omega_{s}(\vec{k})=c_{s}|\vec{k}|-i\Gamma\vec{k}^{2}+\mathcal{O}(\vec{k}^{4})$ Edalati et al. (2010a); Davison and Kaplis (2011); Moitra et al. (2021).
@@
The stress-tensor correlators are computed from the fluctuation equations for the appropriate modes of the metric $g_{MN}$, which couple to the fluctuations of the gauge field (50). These equations can be written in terms of two gauge-invariants Kodama and Ishibashi (2004); Edalati et al. (2010a)
@@
The fields $\mathcal{A}$ and $\Phi$ obey Kodama and Ishibashi (2004); Edalati et al. (2010a)
@@
It is possible to further simplify the problem by introducing the following “master” variables [^14] Note the relations between our variables and those of Kodama-Ishibashi (KI) Kodama and Ishibashi (2004) $\mathcal{A}=\vec{k}^{2}\mathcal{A}_{\text{KI}}\quad,\quad\Phi=\frac{1}{n}\vec{k}^{2}\Phi_{\text{KI}}\,$, $\Phi_{+}=\frac{\vec{k}^{2}}{4nQ}\Phi_{-,\text{KI}}\quad,\quad\Phi_{-}=-\frac{\vec{k}^{2}}{(n+1)(M+\rho)}\Phi_{+,\text{KI}}\,.$
@@
We would now want to relate the stress-tensor and current correlators to the master fields Green’s functions $G^{\pm}(\omega,\vec{k})$. This relation depends on the dimensionality $n$, so we will now consider a fixed value of $n$. The case of $n=2$ was treated in Edalati et al. (2010a), where the sound channel correlators were found to be non-linearly related to $G^{\pm}$, as reviewed in appendix F. Here we will instead consider $n=3$, for which the relation becomes linear. We argue in appendix F that this also holds for any $n\geq 4$.
@@
From the definitions (56), and using the near-boundary equations of motion [^16] The latter imply in particular the Ward identities obeyed by the stress-tensor and current correlators (68)-(70). See Edalati et al. (2010a) for more details of this calculation in the case $n=2$., the coefficients $\Phi_{\pm}^{(0)}$ and $\Phi_{\pm}^{(1)}$ can now be related to the metric and gauge field expansions (66)-(67) as
@@
Now taking the imaginary parts on both sides of equations (77)-(82) - and remembering that the contact terms are real - the sound channel two-sided correlators are found to be linearly related to the master correlators $G^{\pm}_{12}$. Since the latter obey the product formula, the stress tensor and current correlators can therefore also be written as a product over their poles. In particular, the $G^{-}$ correlator has a diffusive pole $\omega_{D}(k)=-iD\vec{k}^{2}+\mathcal{O}(\vec{k}^{4})$, whereas $G^{+}$ features a sound pole $\omega_{s}(\vec{k})=c_{s}|\vec{k}|-i\Gamma\vec{k}^{2}+\mathcal{O}(\vec{k}^{4})$ Arean et al. (2021), as can be seen explicitly in the numerical results presented in appendix G. These results also indicate that the soft poles approach the IR poles in the near-extremal limit, as in (20). Following the discussion of section 2.3, in the near-extremal hydrodynamic regime we may therefore write the $G^{\pm}$ correlators as
$G^{-}_{12}(\omega,\vec{k})=\frac{\gamma_{-}(\omega,\vec{k})\mathcal{G}^{-}_{12}(\omega,\vec{k})}{\omega^{2}-\omega_{D}(\vec{k})^{2}}\left(1+\mathcal{O}(r_{e}T)\right)\,$, (87)
$G^{+}_{12}(\omega,\vec{k})=\frac{\gamma_{+}(\omega,\vec{k})\mathcal{G}^{+}_{12}(\omega,\vec{k})}{(\omega^{2}-\omega_{s}(\vec{k})^{2})(\omega^{2}-\omega_{s}^{*}(\vec{k})^{2})}\left(1+\mathcal{O}(r_{e}T)\right)\,$, (88)
with $\gamma_{\pm}(\omega,\vec{k})$ analytic and free of zeroes (except potentially for $\vec{k}^{2}<0$), and $\mathcal{G}^{\pm}_{12}(\omega,\vec{k})$ the IR CFT1 correlators (44), with conformal dimensions
@@
$\text{Im}G_{\text{L}}(\omega,\vec{k})=\frac{\gamma_{L}(\omega,\vec{k})\text{Im}\mathcal{G}^{+}_{R}(\omega,\vec{k})}{(\omega^{2}-\omega_{D}(\vec{k})^{2})(\omega^{2}-\omega_{s}(\vec{k})^{2})(\omega^{2}-\omega_{s}^{*}(\vec{k})^{2})}\left(1+\mathcal{O}(r_{e}T)\right)\,,\quad\text{L}\in\{T,J,\text{mix}\}\,$, (90)
@@
  • As expected Moitra et al. (2021), the leading order extremal hydrodynamic expressions
@@
    agree with the extrapolation of leading order hydrodynamics to zero temperature Kovtun (2012), up to the modification of the low frequency power-law behavior from $\omega$ to $\omega^{2\Delta_{+}(\vec{k})-1}$;
@@
  • For the energy-energy, density-density and energy-density spectral functions, the height of the sound peak is of order $\mathcal{O}((r_{e}|\vec{k}|)^{2\Delta_{+}(\vec{k})-3})$, whereas the diffusive peak is of order $\mathcal{O}((r_{e}|\vec{k}|)^{4\Delta_{+}(\vec{k})-2})$. This indicates that the diffusive peak is suppressed at low temperatures, as was observed in Davison and Kaplis (2011) for $n=2$, and further provides the precise momentum scaling of this suppression in the extremal hydrodynamic regime. In particular, this scaling indicates that the diffusive peak tends to become more important as the momentum is increased, as was also observed in Davison and Kaplis (2011).
@@ 3.3 Magnetized branes
For our last example, we consider a black brane background with $p=3$, meaning that in the near-extremal regime the theory flows to a CFT2 in the IR. A solution of this sort was found in D’Hoker and Kraus (2009), corresponding to a magnetized black brane at zero density. The action for the system is Einstein-Maxwell
@@
The near-extremal magnetic brane solution D’Hoker and Kraus (2009) interpolates between the BTZ solution (101) near the horizon, and $AdS_{d+1}$ in the UV. For the fields of the ansatz (97), this means that when the boundary is approached ($r\to 0$) the fields behave as
$A(r)=-\log{\left(\frac{r}{\ell}\right)}+\mathcal{O}(r^{d})\quad,\quad f(r)=1+\mathcal{O}(r^{d})\quad,\quad h(r)=1+\mathcal{O}(r^{d})\,$, (104)
@@
The potential (113) is regular everywhere and has UV asymptotics of the form (8) with $\nu^{2}=(d-1)^{2}/4\geq 0$, which ensures that the scalar two-sided correlator $G_{12}(\omega,k_{x},k_{y})$ associated with $\phi$ obeys the product formula without zeroes. The poles of the retarded correlator associated with $\phi$ can be computed from a numerical calculation, whose results are presented in appendix H. These results indicate that $\phi$ does not feature any gapless pole, and its soft poles approach the IR CFT2 poles at low temperature. According to the discussion of section 2.3, the near-extremal limit of $G_{12}$ may then be written as
$G_{12}(\omega,k_{x},k_{y})=\gamma_{e}(\omega,k_{x},k_{y})\mathcal{G}^{k_{y}}_{12}(\omega,k_{x})\left(1+\mathcal{O}(r_{e}T)\right)\,$, (114)
@@
The zero-temperature effective susceptibility $\tilde{\gamma}_{e}$ is a function of $r_{e}^{2}(\omega^{2}-k_{x}^{2})$ and $r_{e}^{2}k_{y}^{2}$, which is free of zeroes and analytic in the regime $r_{e}^{2}(\omega^{2}-k_{x}^{2}),r_{e}^{2}k_{y}^{2}\ll 1$.